# Data driven discovery of human mobility models

### v1

@@ Methods
Inspired by statistical mechanics, Wilson applied the concept of entropy on origin-destination flow data, and deduced the gravity model using the maximum entropy principle[36]. Let $F_{ij}$ denote the flow from place $i$ to $j$, the entropy of a flow distribution $\{F_{ij}\}$ is defined as
$S=-\sum_{i,j}p_{ij}\ln p_{ij}$ (4)
@@
$F_{ij}=A_{i}B_{j}O_{i}D_{j}e^{-\beta c_{ij}}$ (8)
where $A_{i}=\left(\sum_{j}B_{j}D_{j}e^{-\beta c_{ij}}\right)^{-1}$, $B_{j}=\left(\sum_{i}A_{i}O_{i}e^{-\beta c_{ij}}\right)^{-1}$. This form suggests that flow volumes follow an exponential decay with travel cost. The decay parameter $\beta$ arises from the unknown total cost $C$, and usually needs real data to be calibrated. The decay function of distance can be deduced by substituting the relationship between travel cost and distance. For example, given $c_{ij}=kd_{ij}$, we get the exponential decay $e^{-k\beta d_{ij}}$; given $c_{ij}=k\ln d_{ij}$, we get the power-law decay $d_{ij}^{-k\beta}$. Therefore, the discovered exponential-power decay $e^{-d_{ij}^{\alpha}}$ indicates a power-law relationship between travel cost and distance $c_{ij}=\frac{1}{\beta}d_{ij}^{\alpha}$.

### v2

@@ Appendix A Maximum entropy approach to the gravity model
Inspired by statistical mechanics, [57] applied the concept of entropy on origin-destination flow data, and deduced the gravity model using the maximum entropy principle. Let $F_{ij}$ denote the flow from place $i$ to $j$, the entropy of a flow distribution $\{F_{ij}\}$ is defined as
$S=-\sum_{i,j}p_{ij}\ln p_{ij}$ (A.1)
@@
$F_{ij}=A_{i}B_{j}O_{i}D_{j}e^{-\beta c_{ij}}$ (A.5)
where $A_{i}=\left(\sum_{j}B_{j}D_{j}e^{-\beta c_{ij}}\right)^{-1}$, $B_{j}=\left(\sum_{i}A_{i}O_{i}e^{-\beta c_{ij}}\right)^{-1}$. This form suggests that flow volumes follow an exponential decay with travel cost. The decay parameter $\beta$ arises from the unknown total cost $C$, and usually needs real data to be calibrated.
The decay function of distance can be deduced by substituting the relationship between travel cost and distance. For example, given $c_{ij}=kd_{ij}$, we get the exponential decay $e^{-k\beta d_{ij}}$; given $c_{ij}=k\ln d_{ij}$, we get the power-law decay $d_{ij}^{-k\beta}$. Therefore, the discovered exponential-power decay $e^{-d_{ij}^{\alpha}}$ indicates a power-law relationship between travel cost and distance $c_{ij}=\frac{1}{\beta}d_{ij}^{\alpha}$.